Herlige Narren AS has identified an investment project with the following cash flows. Assume all the cash flows have occurred at the beginning of the year. (Do not round intermediate steps. Round your answer to 2 decimal places. The program includes a margin of error of +/-1%.) Year 2 4 Cash Flow 960 1,190 1,540 1,900 Requirement 1: If the discount rate is 9 per cent, what is the future value of these cash flows at the end of year 5? Future value Requirement 2: What is the future value of these cash flows at the end of year 5 at a discount rate of 14 per cent? Future value Requirement 3: What is the future value of these cash flows at the end of year 5 at a discount rate of 20 per cent? Future value
Answer #1
requirement 1
Year Cash flow Future value at the end of year 5
1 960                            1,477.08 =960*1.09*1.09*1.09*1.09*1.09
2 1190                            1,679.78 =1190*1.09*1.09*1.09*1.09
3 1540                            1,994.34 =1540*1.09*1.09*1.09
4 1900                            2,257.39 =1900*1.09*1.09
                           7,408.60
therefore future value =                                                 7,408.60
requirement 2
Year Cash flow Future value at the end of year 5
1 960                            1,848.40 =960*1.14*1.14*1.14*1.14*1.14
2 1190                            2,009.86 =1190*1.14*1.14*1.14*1.14
3 1540                            2,281.58 =1540*1.14*1.14*1.14
4 1900                            2,469.24 =1900*1.14*1.14
                           8,609.08
therefore future value =                                                 8,609.08
requirement 3
Year Cash flow Future value at the end of year 5
1 960                            2,388.79 =960*1.2*1.2*1.2*1.2*1.2
2 1190                            2,467.58 =1190*1.2*1.2*1.2*1.2
3 1540                            2,661.12 =1540*1.2*1.2*1.2
4 1900                            2,736.00 =1900*1.2*1.2
                         10,253.49
therefore future value =                                               10,253.49